You invested $7000 in two accounts paying 8% and 6% annual interest, respectively. If the total interest earned for the year was $448, how much was invested at each rate?
Answer by jorel1380(3719)

.08n+.06(7000-n)=448
8n+6(7000-n)=44800
8n+42000-6n=44800
2n=2800
n=1400
7000-n=5600
$1400 was invested at 8%, and $5600 was invested at 6%..